How to solve for x?
10.2=12*x+7*(1-x)

Answer by colliefan(55)

10. 2=12*x+7*(1-x)
10. 2 = 12*x + 7 - 7*x
10. 2 = 12x -7x +7
10. 2 = 5x +7
10. 2 - 7 = 5x +7 - 7
10. -5 = 5x + 0
10. -5 = 5x
10. -5 * 1/5 = 5x * 1/5
10. -1 = x *5 * 1/5
10. -1 = x * 1
10. -1 = x or x = -1
Be sure to substitute 1 back into the original equation to confirm that the solution is -1.